1) Suppose a perfectly competitive (price-taking) market is made up of 100 identical firms. Each firm...

1)

Suppose a perfectly competitive (price-taking) market is made up of 100 identical firms. Each firm individually has the following costs:

Quantity 0 1 2 3 4 5 6
Total Cost 10 16 20 27 36 46 58

          

Also assume the market demand curve is:

Q Demanded 600 500 400 300 200 100 0
Price 4 7 9 10 12 15 20

            

What is the equilibrium price in this market?

Group of answer choices

a $12

b $9

c $7

d $10
Answer #1

Magiul cost (MC) dCTC) ta (Q) o (16-10] -0) ༧ ) =6 (20-16) = 4 from the the above fique so equilibuumpuice Quantity Toful Cos